Let A, B, and C be independent random variables, uniformly distributed over [0,9], [0,2], and [0,3] respectively. What is the probability that both roots of the equation Ax^2+Bx+C=0 are real?
Answer #1
f(a,b,c)= (1/((a2-a1)*(b2-b1)*(c2-c1))=    1/54
probability=∫c1b2^2/4a2a1a2√4acb2 f(a,b,c) db da dc +∫b2^2/4a2c2b1b20b^2/4c f(a,b,c) da db dc
=f(a,b,c)*(b2*(a2-a1)*(b22/4a2-c1)-(8/9)*((b22/4a2-c1)*(a2-a1))3/2)+f(x)*(b23/12)*(ln(c2)-ln(b22/4a2))
=0.0206+0.0407
=0.0613